ganeshie8 (ganeshie8):

show that \(\large (p-1)!^{p} + 1\) is divisible by \(\large p^2\) \(p\) is an odd prime

OpenStudy (anonymous):

Uh, do you really need help or are you just testing us ?

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

Welcome to OS ^^^

OpenStudy (anonymous):

naa testing the ghost who comes here ;)

OpenStudy (anonymous):

thank, what do I do first?

OpenStudy (turingtest):

I come to these problems so I can feel stupid while I watch @Marki and @ganeshie8 do things I can barely understand. I enjoy the show, though :)

OpenStudy (anonymous):

yaaaaaaaaaay

ganeshie8 (ganeshie8):

XD found this cool property in the morning while trying @PFEH.1999 's recent problem

OpenStudy (vishweshshrimali5):

p is an odd prime. Right ?

ganeshie8 (ganeshie8):

@ReadytoLearn_Student double click below for a spoiler :) \(\color{white}{wilson theorem }\)

ganeshie8 (ganeshie8):

yes p is odd prime, let me add that to the main question thanks vishy :)

OpenStudy (vishweshshrimali5):

:) yw

OpenStudy (vishweshshrimali5):

Okay lets see 1. Parity of both LHS and RHS is same (both are odd) :D

OpenStudy (anonymous):

well idk welson said this \((p-1)!=-1 \mod p\) thus \((p-1)!=-1 \mod p^2\) since p is odd \((p-1)!^p=-1 \mod p^2\) and we get \((p-1)!^p=-1 \mod p^2 \)

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\dfrac{((p-1)!)^p+1}{p^2}\hspace{.33em}\\~\\ \implies \dfrac{1}{p}\times \dfrac{((p-1)!)^p+1}{p}\hspace{.33em}\\~\\ \implies \dfrac{1}{p}\times \dfrac{((-1))^p+1}{p}------(wilson ~theorm)\hspace{.33em}\\~\\ \implies \dfrac{1}{p}\times \dfrac{(-1)+1}{p}------(as~p~is~odd)\hspace{.33em}\\~\\ \implies \dfrac{1}{p}\times \dfrac{0}{p}\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

but 1/p is not integer so ...

ganeshie8 (ganeshie8):

its a very good try, but it breaks in the end p|a need not imply p^n | a

OpenStudy (mathmath333):

yes thats correct

ganeshie8 (ganeshie8):

suppose a|p, below logic wont work : \[\dfrac{a}{p^n} =\dfrac{1}{p^{n-1}}\cdot \dfrac{a}{p}~ \stackrel{\color{red}{?}}{=}~\dfrac{1}{p^{n-1}}\cdot 0 \]

OpenStudy (mathmath333):

but is it true for \((mod ~p)\)

ganeshie8 (ganeshie8):

yes wilson gives us \[(p-1)! \equiv -1 \pmod p\]

OpenStudy (anonymous):

ok what about this \(p(p-1)!=-p \mod p^2\) \((p(p-1)!)^p=-p^p \mod p^2\) \(p^p(p-1)!^p=-p^p \mod p^2\) \(p^p(p-1)!^p+p^p=0 \mod p^2\) \(p^p((p-1)!^p+1)=0 \mod p^2\)

OpenStudy (anonymous):

lol that leads to nothing

ganeshie8 (ganeshie8):

yes you're not allowed to cancel p^p

OpenStudy (anonymous):

i dint cancel i moved it to the other side

ganeshie8 (ganeshie8):

right, im talking about the next step

ganeshie8 (ganeshie8):

to conclude, you need to cancel p^p but you're not allowed to cancel p^p because gcd(p^2, p^p) is not 1

OpenStudy (anonymous):

p^2|p^p so thats why close road

ganeshie8 (ganeshie8):

Exactly ^

OpenStudy (anonymous):

yeah thats why im trying something else without wilson :P

OpenStudy (anonymous):

i give up , i'll just set and see

OpenStudy (mathmath333):

is this an counter example https://www.wolframalpha.com/input/?i=is+%28%2819-1%29%21%29%5E%7B19%7D%2B1+mod+19%5E2%3D0+%3F

OpenStudy (mathmath333):

O_0

OpenStudy (anonymous):

hmm yeah it works now

ganeshie8 (ganeshie8):

Looks perfect!

OpenStudy (anonymous):

:\

ganeshie8 (ganeshie8):

there should be another way to work this w/o using binomial thm euler generalization mightalso work

OpenStudy (anonymous):

we need binomial to generate the (p,p-1)pk term hmm

ganeshie8 (ganeshie8):

thats a good catch :) yes that looks like a typo

OpenStudy (anonymous):

idk how u'll do it with euler (p-1)!^(p^2-p)=1 mod p^2 hmm

OpenStudy (anonymous):

\(\Large \text{from Wilson we got }\\ \Large (p-1)! \equiv -1 \pmod p \text{ so } (p-1)!=pk-1\\ \Large \text{ with binomial we got }\\ \Large \begin{align*} (pk-1)^p&= \sum _{i=0}^p\binom{p}{i}(pk)^{p-i}(-1)^i\\ &= p^2 \sum _{i=0}^{p-2}\binom{p}{i}(pk)^{p-i-2}(-1)^i+\binom{p}{p-1}pk-1\\ &= p^2 \sum _{i=0}^{p-2}\binom{p}{i}(pk)^{p-i-2}(-1)^i+p^2k-1\\ &= p^2 (\sum _{i=0}^{p-2}\binom{p}{i}(pk)^{p-i-2}(-1)^i+k)-1\\ \end{align*}\\\) \(\Large \begin{align*} (p-1)!^{p} &= p^2 (\sum _{i=0}^{p-2}\binom{p}{i}(pk)^{p-i-2}(-1)^i+k)-1 \\ &\equiv p^2 (\sum _{i=0}^{p-2}\binom{p}{i}(pk)^{p-i-2}(-1)^i+k)-1 \mod p^2 \\ &\equiv 0-1 \mod p^2 \end{align*} \)

OpenStudy (anonymous):

i hope no typos left :(

OpenStudy (anonymous):

hmm , idk

OpenStudy (anonymous):

what other method could works

ganeshie8 (ganeshie8):

looks neat to me

OpenStudy (anonymous):

yeah cuz its ur method lol

ganeshie8 (ganeshie8):

its @PFEH.1999 's method

OpenStudy (anonymous):

ohh

ganeshie8 (ganeshie8):

actually euler generalization can be derived from fermat's thm using this method wana try ?

ganeshie8 (ganeshie8):

prove \(\large a^{\phi(n)} \equiv 1 \pmod n\) using Fermat's little theorem and binomial theorem

ganeshie8 (ganeshie8):

try it when free :)

OpenStudy (anonymous):

ohk

OpenStudy (anonymous):

@ganeshie8 , the proof of euler thm using multiplicative group of integers modulo n is more exciting!

ganeshie8 (ganeshie8):

Interesting looks there are many proofs for euler thm xD does that use the fact that phi is a multiplicative function ?

OpenStudy (anonymous):

if ur interesting in knowing the proof

OpenStudy (anonymous):

it depend on the facts, that u can arrange relatively prime in multiple pairs to get order of congruent and gcd(n,all numbers relatively prime with n )=1